Jada3565:

Find the solution for this system of equations. 12x + 15y = 34 -6x + 5y = 3

jhonyy9:

12x +15y = 34 -6x +5y = 3 this is the system of equation so how you need to solve it using what method ?

jhonyy9:

elimination or substitution method ?